The area of a triangular piece of Stainglass is 50 cm². If the height of the triangle is four times the base how long are the height and base of the piece of stained glass?
Answer by addingup(3677)

1/2bh = 50
h= 4b
1/2b(4b)= 50
2b^2 = 50
b^2 = 25
b = 5 and h:
h = 4b = 4(5)= 20
The base is 5 and the height is 20.